OpenStudy (penguin1234):

From a standard deck of cards, what is the probability that you draw a red card, replace it and then draw a heart?

OpenStudy (penguin1234):

Someone please help

OpenStudy (cydney_morgan):

Here, I don't know if this helps much, But I found this answer offline: In a deck there a 52 cards with 4 suits(hearts,spades,diamonds and clovers). 2 suits are red while the other 2 are black. And there are 4 cards for each number.(ace,jack.queen etc.)And remember an ace isnt considered a face card. 1) there are 2 red kings in any given deck so probability is 2/52=1/26 2)since there are 4 cards of each you multiply 4*3=12 probability is 12/52=3/13 3)I have no idea what a fade is. 4)Remember that there are 4 suits? Each suit has 13 cards. So as there are 2 suits probability is 13*2/52=26/52=1/2 5)There are 26 black cards,and they contain 6 face cards. And there are 26 red cards which contain only 2 red aces. so add both 8/52=2/13 6)In a deck there are 12 face cards (king, queen, and jack) and 13 cards of spades 3 face cards are already in spades so subtract 3. (9+13)/52=22/52=11/26

OpenStudy (penguin1234):

I only have answer choices of 1/8 or 3/8 or 3/4 or 2

OpenStudy (cydney_morgan):

.... Err.. @iambatman do you know how to do this?

OpenStudy (anonymous):

Do I have to read all that :c

OpenStudy (cydney_morgan):

No, No you don't batman.

OpenStudy (penguin1234):

please someone help me

jimthompson5910 (jim_thompson5910):

There are 26 red cards (13 heart + 13 diamonds) out of 52 total. So P(Red) = 26/52 = 1/2 ------------------------------------------------------- There are 13 heart cards (A,2-10, J,Q,K) out of 52 total P(Heart) = 13/52 = 1/4 ------------------------------------------------------- Since each event is independent (the card drawn first is replaced), this means P(Red & Heart) = P(Red)*P(Heart) P(Red & Heart) = (1/2)*(1/4) P(Red & Heart) = (1*1)/(2*4) P(Red & Heart) = 1/8